Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 620     Entry time: 16 Jun 2025, 09:41
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 16.06.2025 
General Announcement:
- A. Galler says a second Cryo issue happened on Saturday but it was possible to recovered by 18:00.
- LIMP-25 is about to start in one week. In case of questions please do not hesitate in contacting Martin Doering and XO.

CTRL:
- New broker version is ready to be installed to cure issues saw in SA2 recently.

DA:
- Nothing major to report from DA.

EEE:
- Issues on Saturday morning (during beam downtime) with PLCs in SCS to be restarted, affecting local vacuum interlocks. 
- Camera issue also appeared at SCS but CTRL was working on it.

VAC:
- Nothing to report.

XPD:
- Nothing to report.

DET:
- Everything is fine.

ITDM: 
- Nothing to report.

XRO:
- Nothing to report.

LAS:
- Quiet week but SA3 PP-laser had not enough triggers. Tests with EEE were done and will be discussed with Desy.

UNSYS:
- Nothing to report.

SPB:
- Nothing to report.

FXE:
- Nothing to report; beam stability seems to be fine by now.

MID:
- All fine at MID.

HED:
- Not present.

SCS:
- Vacuum loop problem seems to be resolved.

SQS:
- Nothing to report.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00